# MicroMoE: Fine-grained Load Balancing for Mixture-of-Experts with Token Scheduling

### v1

@@ Abstract.
Mixture-of-Experts (MoE) has emerged as a promising approach to scale up deep learning models due to its significant reduction in computational resources. However, the dynamic nature of MoE leads to load imbalance among experts, severely impacting training efficiency. While previous research has attempted to address the load balancing challenge, existing solutions either compromise model accuracy or introduce additional system overhead. As a result, they fail to achieve fine-grained load balancing, which is crucial to optimizing training efficiency.
@@
We propose MicroEP, a novel parallelization strategy to achieve fine-grained load balancing in MoE systems. MicroEP is capable of achieving optimal load balancing in every micro-batch through efficient token scheduling across GPUs. Furthermore, we propose MicroMoE, an efficient distributed MoE training system with MicroEP’s load balancing capabilities. Our experimental results demonstrate that MicroMoE improves the end-to-end training throughput by up to 47.6% compared with the state-of-the-art system, and almost consistently achieves optimal load balance among GPUs.
distributed training, mixture-of-experts, load balance
1. Introduction
@@
In recent years, transformer-based large models have fueled significant advancements across various domains, including natural language processing (Transformer; GPT-2; PaLM), computer vision (V-MoE; Swin_Transformer), and multimodality (ViLBERT; Cao_2023_ICCV). These achievements underscore that scaling up model size is a straightforward yet fundamental approach to enhancing model performance. Nevertheless, deploying large models necessitates substantial computational resources. To tackle this, Mixture-of-Experts (MoE) proposes a sparsely-activated architecture (shazeer2017outrageouslylargeneuralnetworks), enabling the scaling up of model capacity with only a sub-linear increase in computational cost. Nowadays, MoE has gained widespread adoption in state-of-the-art models, such as GPT-4 (GPT-4), Llama 4 (Llama-4), Gemini 2.5 (Gemeni-2.5), Grok-4 (Grok-4), Phi-3.5-MoE (Phi-3), Claude 4 (Claude4), Qwen3 (Qwen3), and DeepSeek-V3 (DeepSeek-V3).
@@
However, the sparse nature of MoE can indeed introduce inefficiencies during training. MoE divides a model layer into multiple experts and dynamically routes each input token to its top-K most suitable experts at runtime. This dynamic routing can result in significant load imbalances among experts (FasterMoE; GShard). Furthermore, recent distributed training frameworks commonly employ expert parallelism (EP) to distribute experts across multiple GPUs (Switch_Transformer). While EP reduces memory consumption per GPU, it introduces a straggler problem: All GPUs must wait for the most heavily loaded GPU in the EP group to complete computation before proceeding, resulting in wasted computational resources and reduced system throughput. Therefore, the load imbalance issue poses a significant challenge in efficiently training MoE models.
@@
To address the load imbalance issue of MoE, many works propose algorithmic solutions by modifying the routing algorithms. These approaches incorporate load-balancing loss (Switch_Transformer; GLaM; auxiliary-loss-free), route tokens to less suitable experts (LocMoE; expert_choice), or simply drop excess tokens to alleviate the burden on heavily loaded experts (Switch_Transformer; GShard; LocMoE). However, these modifications to model logic can degrade model accuracy and potentially impact model convergence (guo2025advancingexpertspecializationbetter; OpenMoE; Prophet). As a result, they introduce a trade-off between model accuracy and system efficiency (Pangu; qiu2025demonsdetailimplementingload).
@@
Other works propose systematic solutions for load balancing in MoE without compromising the model accuracy. A series of works balance the GPU loads by adjusting the expert-to-GPU mapping (SmartMoE; FlexMoE; Hecate; Pro-Prophet; Prophet; SwiftMoE). For instance, FlexMoE (FlexMoE) adaptively replicates experts based on their popularity, distributing loads of more popular experts to more GPUs. However, they inevitably introduce substantial system overheads when adjusting expert placements. As a result, they introduce a trade-off between expert adjustment overhead and load balancing performance.
The aforementioned trade-offs prevent existing solutions from effectively achieving ideal load balancing, which should satisfy two requirements:
@@
R1: Maximizing system efficiency requires fine-grained load balance among all GPUs. Giant model sizes demand numerous GPUs for training (Pangu; DeepSeek-V3), making even minor load imbalances costly in terms of wasted GPU time. Some algorithmic solutions can achieve perfect load balance by restricting the number of tokens to each expert (GShard). However, their accuracy-balance trade-off necessitates significant compromise of model accuracy to equalize all GPU loads (expert_choice; Megablocks). Many systematic solutions schedule the expert layout to optimize load balancing. Typically, they treat each expert replica as a scheduling unit for communication efficiency (SmartMoE; FlexMoE; Hecate; Prophet). However, scheduling at the expert granularity results in a limited, discrete scheduling space, making it challenging to achieve optimal load balance across a vast, near-continuous space of possible load distributions.
@@
R2: Dynamic expert load distributions require micro-batch-level adaptation at runtime. Data heterogeneity across micro-batches causes severe variations in expert loads, necessitating dynamic load balancing to optimize GPU utilization consistently. Algorithmic solutions using load-balancing loss rely on per-iteration parameter updates for adaptive adjustments (Switch_Transformer; auxiliary-loss-free), which are too slow to handle per-micro-batch load imbalances. Existing systematic solutions involve migrating expert parameters to align with the changing load distributions (FlexMoE; Prophet; SmartMoE). However, the substantial size of these parameters presents a significant challenge to making frequent adjustments seamlessly.
@@
To meet these requirements, we adopt a novel systematic approach by scheduling input tokens to perform load balancing for MoE systems, which has two advantages: First, token scheduling allows a precise control over computational workloads. Unlike expert scheduling solutions, where all replicas of an expert have even loads (FlexMoE; Prophet), token scheduling may allocate uneven loads to each expert replica, thus achieving more fine-grained load balance among GPUs (R1). Second, token scheduling leverages the pathways of existing EP communication. This minimal communication overhead allows efficient load balancing within each micro-batch (R2).
@@
However, performing token scheduling in existing EP frameworks faces two challenges. C1: Current frameworks dispatch tokens within an EP group, which contains exactly one replica of each expert (Megatron-LM; DeepSpeed). As a result, the GPU loads are fixed by expert loads, which are determined by the routing algorithm. To tackle this, we schedule tokens among expert replicas in multiple EP groups. C2: Identical expert placement across EP groups severely limits the scheduling space (see Figure 3(b)). Since a popular expert has identical EP ranks in all EP groups, naïve token scheduling cannot balance loads across different EP ranks, resulting in only local load balance. To tackle this, we tailor the expert placement to expand the scheduling space and achieve global load balance. Furthermore, we theoretically analyze the influence of expert placement on token scheduling and derive the properties of optimal expert placement with graph theory6.1).
@@
Formally, we propose MicroEP, a novel EP strategy that leverages token scheduling for fine-grained load balancing across GPUs. MicroEP optimizes load balancing from two perspectives: (1) From a short-term perspective, MicroEP schedules tokens within every micro-batch using linear programming to minimize the maximum GPU loads (§5.1). To optimize system efficiency, MicroEP reduces the communication volume through locality-aware routing5.2) and minimizes the scheduling latency through efficient overlapping5.4). (2) From a long-term perspective, MicroEP tailors expert placement to optimize the token scheduling performance and mitigate coarse-grained imbalances. Based on our theoretical analysis, MicroEP adopts multiple expert placement strategies (§6.2-6.3) and an adaptive replacement mechanism (§6.4). These optimizations enable MicroEP to maintain complete load balance under dynamic and highly skewed expert loads.
@@
Based on MicroEP, we design MicroMoE, an efficient distributed MoE training system. We implement MicroMoE upon Megatron-LM (Megatron-LM) and evaluate its load balancing performance. Experimental results show that MicroMoE achieves up to 47.6% improvement on end-to-end throughput compared with Megatron-LM and maintains complete load balance even under highly imbalanced workloads.
@@
  • We propose MicroEP to achieve dynamic and fine-grained load balancing among GPUs with token scheduling (§4).
@@
  • We propose several expert placement strategies and an adaptive replacement mechanism to optimize MicroEP’s load balancing capacity (§6).
  • We implement MicroMoE, an efficient distributed MoE training system based on MicroEP. We conduct comprehensive experiments to demonstrate that MicroMoE exhibits superior system efficiency and load balancing capability (§7).
2. Background
2.1. Transformer and MoE
(a) A dense transformer block.
(b) An MoE transformer block.
(c) An MoE transformer block with expert parallelism.
Figure 1. An example of transformer, MoE, and expert parallelism.
Transformer is the state-of-the-art architecture for deep learning models (Transformer). A transformer model consists of multiple sequential blocks, each comprising an attention layer and a feed-forward network (FFN). Figure 1(a) illustrates the architecture of a standard (dense) transformer block.
Figure 1(b) illustrates an MoE transformer block (Switch_Transformer). MoE replaces the FFN with multiple expert FFNs, coordinated by a gate network. After the attention computation, the gate network assigns each token to its top-K most suitable experts. Then, these selected experts process the token and aggregate their results to produce the final result.
2.2. Expert Parallelism and Expert Data Parallelism
The increasing scale of deep learning models has necessitated the development of various parallelization strategies for distributed training across multiple computational devices. We introduce some of them in this section.
Data parallelism (DP) is a fundamental parallelization strategy that partitions input data across devices to enhance training throughput. In the forward pass, each DP rank (e.g., a device) maintains a complete copy of the model parameters while processing distinct micro-batches of data independently. In the backward pass, all devices in a DP group aggregate their gradients for parameter synchronization. Recent advances, such as ZeRO (ZeRO), have further optimized DP’s memory efficiency by eliminating memory redundancies across devices.
@@ 2.2. Expert Parallelism and Expert Data Parallelism
Pipeline parallelism (PP) addresses memory constraints by partitioning model layers (PipeDream; GPipe). While PP introduces lower communication overhead compared with TP, it wastes some computational resources due to pipeline bubbles (GPipe; Zero_bubble).
Expert parallelism (EP) is a widely employed technique in distributed MoE systems (Switch_Transformer). EP is a specialized hybrid parallelization strategy combining elements of DP and TP. By exploiting the inherent sparsity of MoE, EP benefits from both DP’s training efficiency and TP’s memory efficiency. As shown in Figure 1(c), EP partitions the FFN layer by distributing distinct experts across multiple devices. At runtime, the attention layer and the gate network utilize conventional DP to process different tokens on separate devices. Next, devices within the EP group perform an all-to-all (A2A) communication operation to dispatch tokens towards their designated expert devices. Then, each device processes its received tokens with its local experts. Finally, the devices perform another all-to-all operation to return tokens to their original devices.
@@ 2.3. Load Imbalances in MoE
The dynamic token-to-expert assignment in MoE systems inherently results in significant load imbalances across experts, presenting a critical performance bottleneck for distributed training. We conduct an experiment to trace the expert load distribution in a training process with model configurations detailed in Tabel 2. As illustrated in the left part of Figure 2, the expert load distribution exhibits substantial variability and skewness, particularly during the initial training iterations.
The imbalance issue severely impacts resource utilization and training throughput in MoE systems with EP. In current MoE implementations, the FFN computation time of a GPU is approximately proportional to the total number of tokens assigned to the experts on this GPU (Megablocks). Since GPUs within an EP group must synchronize through all-to-all communication both before and after FFN computations, all GPUs must wait until the slowest GPU completes computation before proceeding. Consequently, the system’s overall throughput is bottlenecked by the most heavily loaded device, namely the straggler. This strict synchronization requirement emphasizes the necessity of effective load balancing strategies to optimize MoE training efficiency.
@@
3. Analysis of Token Scheduling
3.1. Benefits of Token Scheduling
Our goal is to achieve fine-grained load balancing for MoE systems within every micro-batch. We aim to find a systematic approach without compromising model accuracy. In MoE, FFN computation requires two key elements: expert parameters and input tokens. Therefore, we can perform load balancing through scheduling either of them.
We adopt token scheduling rather than expert parameter scheduling for two key reasons: (1) Token scheduling provides fine-grained control of GPU loads, which is necessary for optimal load balance (R1). In contrast, expert scheduling only provides coarse-grained load balancing due to the significantly smaller number of experts compared with tokens. (2) Token scheduling incurs minimal overhead, enabling efficient scheduling within every micro-batch (R2). Since EP already involves token dispatching, we can piggyback token arrangements on existing all-to-all operations without extra communication overhead. In contrast, expert scheduling requires costly parameter migration across devices, making it inefficient to execute in each micro-batch. These advantages make tokens, rather than expert parameters, a more practical target for scheduling optimization.
3.2. Challenges in Token Scheduling
@@
(c) MicroEP, shuffling expert placement and scheduling replica loads. EDP groups become {0,3}, {0,1}, {1,2}, {2,3}. Afterward, all GPU loads are equal.
Figure 3. Converting EP to MicroEP. The shape and color of a symbol indicate the source GPU and the assigned expert of a token. The bottom curves indicate EDP groups.
We adopt a token scheduling approach to balance loads across GPUs. However, this approach poses two fundamental challenges:
@@
We aim to balance GPU loads by scheduling input tokens. To achieve this, we must determine the scheduling space. Specifically, when a token is assigned to an expert, we must be able to compute it on one of multiple GPUs.
Unfortunately, the scheduling space is unattainable in the current EP paradigm. As shown in Figure 3(a), vanilla EP restricts token dispatching within the scope of an EP group, where each EP group contains exactly one replica of each expert’s parameters. When a token is assigned to an expert, it must be computed on the GPU hosting that expert’s replica in its EP group. Consequently, the token-to-GPU mapping is fixed by the token-to-expert mapping, eliminating any space for computation scheduling.
@@
Solution 2: To expand the scheduling space, we shuffle the expert placement within the merged EP groups. This operation creates intersecting EDP groups across different experts, substantially enlarging the scheduling space. As illustrated in Figure 3(c), this approach can potentially achieve perfect load balance across all GPUs.
4. MicroMoE Overview
Figure 4. MicroMoE architecture.

### v2

# FineMoE: Fine-grained Load Balancing for Mixture-of-Experts with Token Scheduling

@@ Abstract
Mixture-of-Experts (MoE) has emerged as a promising approach to scale up deep learning models due to its significant reduction in computational resources. However, the dynamic nature of MoE leads to load imbalance among experts, severely impacting training efficiency. While previous research has attempted to address the load balancing challenge, existing solutions either compromise model accuracy or introduce additional system overhead. As a result, they fail to achieve efficient and fine-grained load balancing, which is crucial to optimizing training efficiency.
@@
We propose FineEP, a novel parallelization strategy to achieve fine-grained load balancing in MoE systems. FineEP is capable of achieving complete load balancing in every micro-batch through efficient token scheduling across GPUs. Furthermore, we propose FineMoE, an efficient distributed MoE training system with FineEP’s load balancing capabilities. Our experimental results demonstrate that FineMoE improves the end-to-end training throughput by up to 47.6% compared with the state-of-the-art system, and almost consistently achieves complete load balance among GPUs.
1 Introduction
@@
In recent years, transformer-based large models have fueled significant advancements across various domains, including natural language processing [Transformer, GPT-2, PaLM], computer vision [V-MoE, Swin_Transformer], and multimodality [ViLBERT, Cao_2023_ICCV]. These achievements underscore that scaling up model size is a straightforward yet fundamental approach to enhancing model performance. Nevertheless, deploying large models necessitates substantial computational resources. To tackle this, Mixture-of-Experts (MoE) proposes a sparsely-activated architecture [shazeer2017outrageouslylargeneuralnetworks], enabling the scaling up of model capacity with only a sub-linear increase in computational cost. Nowadays, MoE has gained widespread adoption in state-of-the-art models, such as GPT-5 [GPT-5], Llama 4 [Llama-4], Gemini 3 [Gemeni-3], Grok-4 [Grok-4], Claude 4 [Claude4], Qwen3 [Qwen3], and DeepSeek-V3 [DeepSeek-V3].
@@
However, the sparse nature of MoE can indeed introduce inefficiencies during training. MoE divides a model layer into multiple experts and routes each input token to its top-K most suitable experts at runtime. This dynamic routing can result in significant load imbalances among experts [FasterMoE, GShard]. Furthermore, recent distributed training frameworks commonly employ expert parallelism (EP) to distribute experts across multiple GPUs [Switch_Transformer]. While EP reduces memory consumption per GPU, it introduces a straggler problem: All GPUs must wait for the most heavily loaded GPU in the EP group to complete computation before proceeding, resulting in wasted computational resources and reduced system throughput. Therefore, the load imbalance issue poses a significant challenge in efficiently training MoE models.
@@
To address the load imbalance issue of MoE, many works propose algorithmic solutions by modifying the token-to-expert routing algorithm. These approaches incorporate load-balancing loss [Switch_Transformer, GLaM, auxiliary-loss-free], route tokens to less suitable experts [LocMoE, expert_choice], or simply drop excess tokens to alleviate the burden on heavily loaded experts [Switch_Transformer, GShard, LocMoE]. However, these modifications to model logic can degrade model accuracy and potentially impact model convergence [guo2025advancingexpertspecializationbetter, OpenMoE, Prophet]. As a result, they introduce a trade-off between model accuracy and system efficiency [Pangu, qiu2025demonsdetailimplementingload].
@@
Other works propose systematic solutions for load balancing in MoE without compromising the model accuracy. A series of works balance GPU loads through expert scheduling, adjusting the expert-to-GPU mapping [SmartMoE, FlexMoE, Hecate, Pro-Prophet, Prophet, SwiftMoE]. For instance, FlexMoE [FlexMoE] adaptively replicates experts based on their popularity, distributing loads of more popular experts to more GPUs. However, these solutions cannot effectively achieve ideal load balance, which should satisfy two fundamental requirements:
@@
R1: Maximizing system efficiency requires fine-grained load balance among all GPUs. Giant model sizes demand numerous GPUs for training [Pangu, DeepSeek-V3], making even minor load imbalances costly in terms of wasted GPU time. The significant expense necessitates a fine granularity for load balancing, hopefully equalizing all GPU loads. However, existing expert scheduling solutions typically treat each expert replica as a scheduling unit for communication efficiency [SmartMoE, FlexMoE, Hecate, Prophet]. Since scheduling at the expert granularity results in a limited, discrete scheduling space, they fail to achieve optimal load balance across a vast, near-continuous space of possible load distributions.
@@
R2: Dynamic expert load distributions require micro-batch-level adaptation at runtime. Data heterogeneity across micro-batches causes severe variations in expert loads, necessitating dynamic load balancing to optimize GPU utilization consistently. Existing expert scheduling solutions involve migrating expert parameters to align with the changing load distributions [FlexMoE, Prophet, SmartMoE]. However, the substantial size of these parameters presents a significant challenge to making frequent adjustments seamlessly.
@@
To meet these requirements, we adopt a novel systematic approach by scheduling input tokens instead of experts, which has two advantages: (R1) First, token scheduling allows a precise control over per-GPU workloads (relative to token counts), enabling fine-grained load balancing. (R2) Second, token scheduling leverages existing all-to-all communication in EP, incurring minimal communication overhead. This allows us to achieve dynamic load balancing within each micro-batch without compromising the entire system efficiency. Furthermore, we can combine fine-grained token scheduling and coarse-grained expert scheduling, improving overall load balancing capability.
@@
However, performing token scheduling in existing EP frameworks faces two challenges. (C1) Current frameworks dispatch tokens within an EP group, which contains exactly one replica of each expert [Megatron-LM, DeepSpeed]. As a result, the GPU loads are fixed by expert loads, which are determined by the routing algorithm. To tackle this, we schedule tokens among expert replicas across multiple EP groups, balancing GPU loads while keeping expert loads. (C2) Identical expert placement across EP groups severely limits the scheduling space (see Figure 3(b)). Since replicas of a popular expert have identical EP ranks in all EP groups, naïve token scheduling cannot migrate their loads to other EP ranks, resulting in only local load balance. Although some studies adopt non-identical expert placement [FlexMoE], the relationship between expert placement and token scheduling remains uninvestigated. To tackle this, we theoretically analyze the influence of expert placement on token scheduling with graph theory6.1). Based on the analysis, we tailor the expert placement to expand the scheduling space and achieve global load balance.
@@
We propose FineEP, a novel EP strategy that leverages token scheduling for fine-grained load balancing across GPUs. FineEP optimizes load balancing from two perspectives: (1) From a short-term perspective, FineEP schedules tokens within every micro-batch using linear programming to minimize the maximum GPU loads (§5.1). To optimize system efficiency, FineEP reduces communication overhead through locality-aware routing5.2) and minimizes scheduling latency through overlapping5.4). (2) From a long-term perspective, FineEP tailors expert placement to optimize its token scheduling performance and mitigate coarse-grained imbalances. Based on our theoretical analysis, FineEP incorporates multiple expert placement strategies (§6.2-6.3) and an adaptive replacement mechanism (§6.4). These optimizations enable FineEP to maintain complete load balance under dynamic and highly skewed workloads.
@@
Based on FineEP, we design FineMoE, an efficient distributed MoE training system. We implement FineMoE upon Megatron-LM [Megatron-LM] and evaluate its load balancing performance. Experimental results show that FineMoE achieves up to 47.6% improvement on end-to-end throughput compared with Megatron-LM and maintains complete load balance even under highly imbalanced workloads.
@@
  • We propose FineEP to achieve dynamic and fine-grained load balancing among GPUs with token scheduling (§4).
@@
  • We propose several expert placement strategies and an adaptive replacement mechanism to optimize FineEP’s load balancing capacity (§6).
  • We implement FineMoE, an efficient distributed MoE training system based on FineEP. We conduct comprehensive experiments to demonstrate that FineMoE exhibits superior system efficiency and load balancing capability (§7).
2 Background
2.1 Transformer and MoE
(a) A dense transformer layer.
(b) An MoE transformer layer.
(c) An MoE transformer layer with expert parallelism.
Figure 1: An example of transformer, MoE, and expert parallelism.
Transformer is the state-of-the-art architecture for deep learning models [Transformer]. A transformer model consists of multiple sequential layers, each comprising an attention layer and a feed-forward network (FFN). Figure 1(a) illustrates the architecture of a standard (dense) transformer layer.
Figure 1(b) illustrates an MoE transformer layer [Switch_Transformer]. MoE replaces the FFN with multiple expert FFNs, coordinated by a gate network. After the attention computation, the gate network assigns each token to its top-K most suitable experts. Then, these selected experts process the token and aggregate their results to produce the final result.
2.2 Expert Parallelism and Expert Data Parallelism
The increasing scale of deep learning models has necessitated the development of various parallelization strategies for distributed training across multiple computational devices. We introduce some of them as follows.
Data parallelism (DP) is a fundamental parallelization strategy that partitions input data across devices to enhance training throughput. In the forward pass, each DP rank (e.g., a device) maintains a complete copy of the model parameters while processing distinct micro-batches of data independently. In the backward pass, all devices in a DP group aggregate their gradients for parameter synchronization. Recent advances, such as ZeRO [ZeRO], have further optimized DP’s memory efficiency by eliminating memory redundancies across devices.
@@ 2.2 Expert Parallelism and Expert Data Parallelism
Pipeline parallelism (PP) addresses memory constraints by partitioning model layers [PipeDream, GPipe]. While PP introduces lower communication overhead compared with TP, it wastes some computational resources due to pipeline bubbles [GPipe, Zero_bubble].
Expert parallelism (EP) is a widely employed technique in distributed MoE systems [Switch_Transformer]. EP is a specialized hybrid parallelization strategy combining elements of DP and TP. By exploiting the inherent sparsity of MoE, EP benefits from both DP’s training efficiency and TP’s memory efficiency. As shown in Figure 1(c), EP partitions the FFN layer by distributing distinct experts across multiple devices. At runtime, the attention layer and the gate network utilize conventional DP to process different tokens on separate devices. Next, devices within the EP group perform an all-to-all communication operation to dispatch tokens towards their designated expert devices. Then, each device processes its received tokens with its local experts. Finally, the devices perform another all-to-all operation to return tokens to their original devices.
@@ 2.3 Load Imbalances in MoE
The dynamic token-to-expert assignment in MoE systems inherently results in significant load imbalances across experts, presenting a critical performance bottleneck for distributed training. We conduct an experiment to trace the expert load distribution in a training process. As illustrated in the left part of Figure 2, the expert load distribution exhibits substantial variability and skewness, particularly during the initial training iterations.
The imbalance issue severely impacts resource utilization and training throughput in MoE systems with EP. In current MoE implementations, the FFN computation time of a GPU is approximately proportional to the total number of tokens assigned to the experts on this GPU [Megablocks]. Since GPUs within an EP group must synchronize through all-to-all communication both before and after FFN computation, all GPUs must wait until the slowest GPU completes computation before proceeding. Consequently, the system’s overall throughput is bottlenecked by the most heavily loaded device, namely the straggler. This strict synchronization requirement emphasizes the necessity of effective load balancing strategies to optimize MoE training efficiency.
@@
3 Challenges in Token Scheduling
@@
(c) FineEP, shuffling expert placement and scheduling replica loads. Afterward, all GPU loads are equal.
Figure 3: Converting EP to FineEP. The shape and color of a symbol indicate the source GPU and the assigned expert of a token. The bottom curves indicate EDP groups.
We aim to achieve fine-grained load balancing for MoE systems through token scheduling within every micro-batch. However, this approach poses two fundamental challenges:
@@
To perform token scheduling for load balancing, we must first determine the scheduling space. Specifically, when a token is assigned to an expert, we must be able to compute it on one of multiple GPUs.
Unfortunately, such scheduling space is unattainable in the vanilla EP paradigm. As shown in Figure 3(a), vanilla EP restricts token dispatching within the scope of an EP group, where each EP group contains exactly one replica of each expert’s parameters. When a token is assigned to an expert, it must be computed on the GPU hosting that expert’s replica in its EP group. Consequently, the token-to-GPU mapping is fixed by the token-to-expert mapping, eliminating any space for computation scheduling.
@@
Solution 2: To expand the scheduling space, we shuffle the expert placement within the merged EP groups. This operation creates intersecting EDP groups across different experts, substantially enlarging the scheduling space. As illustrated in Figure 3(c), this approach can potentially achieve complete load balance across all GPUs.
4 FineMoE Overview
Figure 4: FineMoE architecture.